OpenStudy (anonymous):

solve equation, x+2/10=3/10, for given variable. write answer as reduced fraction.

OpenStudy (anonymous):

if \[\frac{x+2}{10}=\frac{3}{10}\]then it means \[x+2=3\]

OpenStudy (anonymous):

1/10?

OpenStudy (anonymous):

no just \(1\)

OpenStudy (anonymous):

if \(x+2=3\) then \(x=1\)

OpenStudy (anonymous):

it was wrong:(

OpenStudy (anonymous):

what was wrong?

OpenStudy (anonymous):

the final answer was supposed to be typed 1/10 wah lol.. online school blows.

OpenStudy (anonymous):

hold on maybe it would be best to have a screen shot of the actual question

OpenStudy (anonymous):

perhaps it was \[x+\frac{2}{10}=\frac{3}{10}\]